OpenStudy (lgbasallote):

what is the corresponding linear DE to the particular solution \[y = 3x^2 - \cosh 2x\] how to do this?

OpenStudy (anonymous):

take derivative and remember you have eliminate the constant here it is 3..

OpenStudy (lgbasallote):

basically i have no idea what root corresponds to cosh or any hyperbolic function

OpenStudy (lgbasallote):

why take the derivtive?

OpenStudy (anonymous):

Don't you want to find DE??

OpenStudy (anonymous):

DE is differential equation.. You are given with solution.. And you have to find corresponding DE.

OpenStudy (wasiqss):

why derivative O.o

OpenStudy (anonymous):

Do you know the form of DE ??

OpenStudy (lgbasallote):

im given a particular solution...i need to find the DE

OpenStudy (lgbasallote):

so wha?

OpenStudy (anonymous):

That is what I am saying.. Take the derivative first time..

OpenStudy (lgbasallote):

so whenever im given a solution and i need a diff eq i take the derivative?

OpenStudy (anonymous):

Yes.. you will..

OpenStudy (lgbasallote):

hmm \[y' = 6x - 2\sin h 2x\] wat now?

OpenStudy (anonymous):

you have to eliminate all the constants..

OpenStudy (anonymous):

Or may be @UnkleRhaukus will help you better.. ha ha ha..

OpenStudy (lgbasallote):

what do you mean??

OpenStudy (anonymous):

Call Uncle Rocks..

OpenStudy (lgbasallote):

lol..you start it and then you let others finish :p that's not fair

OpenStudy (anonymous):

Sorry I don't want to give you false or wrong knowledge..

OpenStudy (anonymous):

Take derivative one more time: \[y'' = 6 - 4\cosh(2x)\]

OpenStudy (anonymous):

By taking derivative one more time 6 will be eliminated: \[y'' = -8\sinh(2x)\]

OpenStudy (lgbasallote):

...and i still dont know the roots that correspond to hyperbolic functions :(

OpenStudy (anonymous):

May Be this is your answer: \[y'' + 8\sinh(2x) = 0\]

OpenStudy (lgbasallote):

actually it's (D^4 - 4D^2); y=0 dunno what that means

OpenStudy (anonymous):

You are given with: \(y = 3x^2 - \cosh 2x\) Can you take it derivative four times ??? Do it hurry up..

OpenStudy (anonymous):

Wait let me try first..

OpenStudy (anonymous):

No I am not getting this actually..

OpenStudy (lgbasallote):

lol =))) hard sint it

OpenStudy (anonymous):

Meaning?

OpenStudy (lgbasallote):

hard isnt it* lol

OpenStudy (anonymous):

No it is not hard.. I studied 4 years before about DE.. My concepts gone far away from me.. Uncle will help you now..

OpenStudy (lgbasallote):

i hope so

OpenStudy (unklerhaukus):

\[\cosh x=\frac{e^x+​e^{​-x}}{​2}\]

OpenStudy (lgbasallote):

oh yeahhh

OpenStudy (lgbasallote):

so cosh2x is....?

OpenStudy (anonymous):

Replace x with 2x..

OpenStudy (lgbasallote):

mmhmm

OpenStudy (lgbasallote):

from there i derive?

OpenStudy (unklerhaukus):

break \(y_p \) up into \(y_{p_1}+y_{p_2}+y_{p_3}\)

OpenStudy (lgbasallote):

that means what again?

OpenStudy (unklerhaukus):

\[y_p = Ax^2 - B\cosh 2x\]\[y_{p_1}=Ax^2\]\[y_{p_2}=Be^{x}\]\[y_{p_3}=Ce^{-x}\]

OpenStudy (lgbasallote):

woah what? how?

OpenStudy (unklerhaukus):

those constant are not ment to be the same

OpenStudy (lgbasallote):

but why? how? wwhat?

OpenStudy (unklerhaukus):

take the derivatives of y_p_1 and find A

OpenStudy (lgbasallote):

why do you have Be^x and Ce^-x

OpenStudy (unklerhaukus):

they came form the cosh x

OpenStudy (lgbasallote):

but it's cosh2x

OpenStudy (anonymous):

A B and C can take any values they are constants.. You have changed cosh(2x) into exponential terms..

OpenStudy (unklerhaukus):

well stick some 2's on

OpenStudy (lgbasallote):

eh?

OpenStudy (unklerhaukus):

in the exponents

OpenStudy (unklerhaukus):

Be^2x and Ce^-2x

OpenStudy (lgbasallote):

what do you mean stick some 2's on?

OpenStudy (lgbasallote):

oh

OpenStudy (unklerhaukus):

like that .

OpenStudy (lgbasallote):

so where did 1/2 go off to?

OpenStudy (unklerhaukus):

it got absorbed in the constant like waterineyes was saying

OpenStudy (lgbasallote):

ohh..okay so what's next?

OpenStudy (unklerhaukus):

\[y^\prime_{p_1}=\]

OpenStudy (lgbasallote):

2A x